. A 9,100 kg tractor is moving (horizontally) at 5.0 km h−1 when the 80 kg driver has
to brake suddenly to come to a stop before hitting a tree stump. If the tractor and
driver come to rest over a horizontal distance of 2.0 m, and the acceleration of the
tractor and driver is constant, what is the magnitude of the net force on the driver
1
Expert's answer
2020-06-10T18:22:20-0400

Initial velocity is "v_0=5.0 km\/h= 1.389 m\/s"

Final velocity is "v=0"

Distance traveled is "d=2.0 m"

The acceleration is

"a=\frac{v^2-v_0^2}{2d}=\frac{(0)^2-(1.389)^2}{2(2.0)}=-0.482 m/s^2"

the magnitude of the net force on the driver is

"F=ma=(80 kg)(0.482 m\/s^2)=38.58 N=38.6 N"